OpenStudy (anonymous):

Write the equation of the line that passes through (3,5) and (6,6) in slope-intercept form. I M P O R T A N T .

OpenStudy (anonymous):

Do you know what form is slope-intercept form?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

do you know how to find slope between two points?

OpenStudy (anonymous):

Yes. This Is How Far I Got...

OpenStudy (anonymous):

y2-y1 = 6-5 =1 x2-x1 - 6-(3) 3

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

So if you have the slope and the form, you've already got the hardest part done. All you have to do now is choose one of your points (say 3,5) and plug in 3 for the x, 5 for the y, and solve for b (the x intercept)

OpenStudy (anonymous):

sorry... *y intercept*

OpenStudy (anonymous):

I'm Still Confused. :\

OpenStudy (anonymous):

Okay, so you start with y=mx+b

OpenStudy (anonymous):

you know that m is (1/3), so now you have y=(1/3)x+b

OpenStudy (anonymous):

Y=(1/3)X+4

OpenStudy (anonymous):

Rakesh Sonai, IKnow That's The Answer But INeed The Work.

OpenStudy (anonymous):

So take your equation and a point (any point; it doesn't matter). Let's do (3,5)

OpenStudy (anonymous):

sorry, just lost internet connection.

OpenStudy (anonymous):

So you know that 3 is your x value and 5 is your y value. Replace every x in your equation with 3 Replace every y in your equation with 5

OpenStudy (anonymous):

Now you have 5=(1/3)(3)+b solve for b

OpenStudy (anonymous):

But Why Am ISupposed To Point It.? First, I'm Supposed To Find The Slop.e Then, I'm Supposed To Choose Either Point And Substitute It Along With The Slope Intercept Form. Last, Solve For B.

OpenStudy (anonymous):

Okay. Here's An Example Of How I'm Supposed To Do It.

OpenStudy (anonymous):

Umme, I like your free-lance use of capital letters :)

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Now Here's The Example. y2-y1=6-5 1 x2-x1=6-(3) 3

OpenStudy (anonymous):

Wait No.

OpenStudy (anonymous):

y2-y1=-1-5 = -6 x2-x1-3-(0) = 3 Slope=-2

OpenStudy (anonymous):

You said the directions were 1) find slope 2)choose point and substitute it, and 3) solve for b. That's what we did. I don't know what you mean by "point it"

OpenStudy (anonymous):

y=mx+b -1=-2(3) -1=-6+6

OpenStudy (anonymous):

You Said Point It. Nevermind.

OpenStudy (anonymous):

So What's The Whole Work.? :\

OpenStudy (anonymous):

Sorry; didn't mean to. Other than the point it thing, are you getting the problem or not really?

OpenStudy (anonymous):

Umme...?

OpenStudy (anonymous):

Yes.IAm.

OpenStudy (anonymous):

The Only Part I'm Confused In Is How To Get The 4. :\

OpenStudy (anonymous):

That's what you get when you solve for b by plugging in the point

OpenStudy (anonymous):

How Though.?

OpenStudy (anonymous):

The Last Thing IGot Was, 6=1/3 (6)+b

OpenStudy (anonymous):

so you put in a 3 where you see x and a 5 where you see y. The only variable you have left is b, so you solve for b. That will give you 4. Here-- let me work it out

OpenStudy (anonymous):

y=(1/3)x+b

OpenStudy (anonymous):

5=(1/3)3+b <---I put in the point for x and y

OpenStudy (anonymous):

5=1+b 4=b

OpenStudy (anonymous):

So now you take the b (4) and put that where you saw b in the equation

OpenStudy (anonymous):

y=(1/3)x+4

OpenStudy (anonymous):

Where Did The 5 Come From.?

OpenStudy (anonymous):

The 5 is the y value. Since the point was (3,5), we know that we'll put a 5 where we see y and there's a y in the beginning of the equation

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

If that's clear, I'm leaving. Hope this helped :)

OpenStudy (anonymous):

 this is the much simplier way to do this. once you findd the slope plug into the formula and put the equation in point slope form or standards